a division wide aptitude test in mathematics was conducted to 1,000 pupils the mean of the test is 58 and the standard deviation is 12. find the z-value that corresponds to a score x = 70. draw the normal curve distribution. what is your conclusion regarding the scores of the pupils

1
Expert's answer
2022-04-26T00:12:16-0400

"Z=\frac{x-\mu}{\sigma}=\frac{70-58}{12}=1"

Conclusion is that 84.13% of pupils have scores less than 70.